# End-to-End Portfolio Optimization with Quantum Annealing

### v1

@@ V Conclusions
This exploration into the integration of quantum annealing with conventional portfolio optimization techniques has revealed substantial potential for enhancing the efficiency and speed of investment strategies. Our research showcases the capability of quantum annealing in addressing complex optimization problems. The hybrid quantum-classical models we have examined exhibit a synergy that harnesses the strengths of both computational approaches, providing a sophisticated toolkit for financial analysts and investors.
As technological advancements continue to reshape the financial industry, the applications of quantum technology are expected to expand and become more sophisticated. The potential of quantum technology to significantly transform financial strategies is immense, with far-reaching implications for risk management, asset allocation, and beyond. Nonetheless, challenges such as quantum hardware development, algorithmic stability, and practical implementation need to be addressed to fully harness this potential.
Future research should focus on improving the scalability of quantum methods, enhancing their integration with classical financial models, and developing more robust quantum infrastructures. As these technologies evolve, they are poised to play a crucial role in redefining the landscape of financial portfolio optimization. The current results pertain to a single individual investor; therefore, future work should explore testing these hybrid and fully quantum methods across multiple users in real-time environments.
Advancing the comprehension and application of quantum technology within the financial sector portends a transformative era in portfolio management characterized by increased efficiency and alignment with the multifaceted and rapid dynamics of global markets. This evolution promises to augment the strategic capabilities of financial institutions, furnishing them with novel methodologies to optimize asset allocation and risk management strategies in an environment marked by heightened market volatility. Consequently, individual investors are anticipated to gain access to sophisticated tools that enhance investment returns while effectively navigating the intricacies of risk management.
Acknowledgement
The authors would like to thank D-Wave Systems for the access to their quantum annealer. We would also like to thank Prof. Greg Byrd for the meaningful discussions and comments.

### v2

@@ V Conclusions
This work presents an end-to-end framework for portfolio optimization that integrates classical mean–variance modeling with quantum annealing-based rebalancing. The experiments demonstrate that, even under current hardware constraints, quantum annealers can successfully handle constrained portfolio formulations and yield allocations comparable to classical baselines.
However, the scope of the present study remains limited to a modest asset universe and should be interpreted as an initial feasibility test rather than a demonstration of large-scale deployment. The results are influenced by inherent limitations of current annealing technology, including restricted problem size due to embedding overhead, limited parameter precision, and the need for hybrid solver orchestration that introduces classical latency. Furthermore, the stochastic nature of the annealing process may lead to solution variability, necessitating multiple runs and statistical averaging.
These constraints imply that the current results reflect the near-term capabilities of quantum annealers rather than their ultimate performance potential. Given the strength of classical MIQP and metaheuristics (GA/SA), our results should be interpreted as evidence of feasibility rather than superiority. Future research will aim to benchmark larger portfolios, explore advanced hybrid workflows, and leverage improvements in annealing hardware and problem embedding to approach industrial-scale portfolio optimization.
Future Work
While this study demonstrates the feasibility of integrating quantum annealing into a practical portfolio construction pipeline, several important avenues remain open for future exploration. The present experiments use a modest asset universe motivated by the constraints of our industry collaboration and the hybrid solver configurations available at the time of execution. Extending the analysis to broader such as the global equity indices, or multi-asset datasets will be essential for assessing scalability and stress-testing the proposed pipeline under realistic institutional workloads. The current quarterly rebalancing procedure can be extended by explicitly modeling transaction costs, slippage, and liquidity constraints. Incorporating these frictions will produce a more realistic assessment of turnover and its impact on realized performance. Given evolving access policies for commercial annealers, future work will explore alternative quantum and quantum-inspired backends, including simulated quantum annealers, Fujitsu DAU, Toshiba SBM, and emerging open-access annealing platforms. This will allow the pipeline to be tested under diverse hardware characteristics and will facilitate reproducibility. Overall, this work provides an initial proof-of-concept integration of quantum annealing into a portfolio-optimization workflow. Future efforts will focus on broadening empirical scope, strengthening baselines, improving reproducibility, and evaluating scalability across markets, datasets, and hardware platforms.
Acknowledgment
The authors would like to thank D-Wave Systems for the access to their quantum annealer. We would also like to thank Prof. Greg Byrd for the meaningful discussions and comments. This work is supported by the NSF, Office of Strategic Initiatives, under Grant No. OSI-2328774 (SM,KI)
Data and code Availability
All code and processed datasets required to reproduce the results will be provided by the authors upon reasonable request.